# Regularization of the ensemble Kalman filter using a non-parametric, non-stationary spatial model

### v3

@@ Appendix H Bandpass filtering of a locally stationary signal
$\frac{\operatorname{\mathbb{E}}|\delta v|}{\operatorname{\mathbb{E}}\utilde{v}}\leq\left(\frac{L_{\cal H}}{\Lambda}\right)^{2}.$ (67)
This equation assesses the mean absolute error in the approximate Eq. (9).
Appendix I Consistency of the estimator of local spectra
Here we show that, with a specific smoothness constraint imposed on the local spectrum ${\bf f}=\{f_{\ell}\}^{\ell_{\rm max}}_{\ell=0}$, the bandpass filters used in the estimator of local spectra (section 3.1) can be selected dependent on the sample size $M$ in such a way that the resulting mean absolute estimation error vanishes as $M\to\infty$.
To define in which sense the local spectrum is smooth, we switch from wavenumbers $\ell$ to log-wavenumbers $\nu=\log(\ell+1)$ and rescale the log-wavenumber coordinate as follows: $\vartheta=\pi\nu/\max(\nu)$ for $\ell=0,1,\dots,\ell_{\rm max}$. Then, we set $\vartheta(-\ell)=\vartheta(\ell)$ so that $\vartheta$ ranges from $-\pi$ to $\pi$. Defining $F(\vartheta(\ell))=f_{\ell}$, we postulate, additionally, that the spatial resolution is large enough to treat $F(\vartheta)$ as a function (of continuous argument) on the unit circle.
Now, we say that the spectrum $f_{\ell}$ is smooth if $F(\vartheta)$ is a band-limited function:
$F(\vartheta)=\sum_{-K+1}^{K}\widetilde{F}_{k}\,\mathrm{e}^{\mathsf{i}k\vartheta},$ (68)
where $\widetilde{F}_{k}$ is the the Fourier transform of $F(\vartheta)$ and $K$ is a measure of smoothness of the spectrum (the lower the smoother).
Smooth spectra and high spatial resolution allow us to replace sums over wavenumbers by integrals. Equation (9) can, then, be rewritten as
$v_{j}=\utilde{v}_{j}+\delta^{\rm nst}_{j},$ (69)
where
@@ Appendix I Consistency of the estimator of local spectra
$T$ is a real number that controls the width of the function $\omega_{j}(\cdot;T)$ (the way it does so is specified below) and $\delta^{\rm nst}_{j}$ is the discrepancy between $\utilde{v}_{j}$ and the true ${v}_{j}$ due to spatial non-stationarity of the random process in question. The discrepancy $\delta^{\rm nst}_{j}$ is studied in H, where it is generically denoted as $\delta v$.
We specify the transfer functions $\omega_{j}(\vartheta)$ of the bandpass filters in Eq. (70) as follows. First, we let the number of the filters, $J$, be equal to $K+1$. Second, we define the function $\Omega:{\mathbb{R}}\to{\mathbb{R}}$ is a positive even function that is monotonic on ${\mathbb{R}}_{+}$, vanishes at infinity, and satisfies for some $\beta>0$
@@
where $\vartheta_{j}=\frac{\pi}{K}(j-1)$, $j=1,2,\dots,J$, determines the $j$th passband’s central log-wavenumber. We note that by definition, Eqs. (72) and (73), $T$ determines the the length scale of the resulting filters’ impulse response functions: $L_{{\cal H}_{j}}=O(T^{-1})$ as $T\to\infty$. As a result, Eq. (67) entails
$\operatorname{\mathbb{E}}|\delta^{\rm nst}_{j}|\leq\left(\frac{L_{{\cal H}_{j}}}{\Lambda}\right)^{2}\operatorname{\mathbb{E}}{\utilde{v}}_{j}=O(T^{-2}).$ (74)
The data we have are the sample band variances $d_{j}=\widehat{v}_{j}$, which estimate $v_{j}$ up to the sampling error $\delta^{\rm smp}_{j}$:
$d_{j}=v_{j}+\delta^{\rm smp}_{j}.$ (75)
@@
$\operatorname{\mathbb{E}}|\delta^{\rm smp}_{j}|\leq\sqrt{\operatorname{\mathbb{E}}(\delta^{\rm smp}_{j})^{2}}\sim\sqrt{\frac{2}{M}}\operatorname{\mathbb{E}}v_{j}=O(M^{-\nicefrac{{1}}{{2}}})$ (76)
as $M\to\infty$. From Eqs. (69), (70), and (75), we have
$\int_{{\mathbb{S}}^{1}}\omega_{j}(\vartheta;T)\,F(\vartheta)\,{\rm d}\vartheta=d_{j}-\delta^{\rm smp}_{j}-\delta^{\rm nst}_{j}.$ (77)
@@
This is a (linear) inverse problem to be solved with respect to $F(\cdot)$ from the data $\{d_{j}\}_{j=1}^{J}$. Due to Eq. (73), the left-hand side of Eq. (77) is the convolution evaluated at $\vartheta_{j}$ of the function $\omega(\vartheta;T)$ with the function $F(\vartheta)$. This convolution can readily be computed in Fourier space, so Eq. (77) becomes
$\widetilde{F}_{k}\,\widetilde{\omega}_{k}\propto\widetilde{d}_{k}-\widetilde{\delta_{k}^{\rm smp}}-\widetilde{\delta^{\rm nst}_{k}},$ (78)
@@
where the variables with the tilde are Fourier transforms of the respective variables without the tilde and $0\leq k\leq K$. From Eq. (78), we can write down the solution in Fourier space, $\widetilde{\widehat{F}}_{k}\propto{\widetilde{d}_{k}}/{\widetilde{\omega}_{k}}$, and its error (up to a constant factor):
$|\delta\widetilde{\widehat{F}}_{k}|\leq\frac{|\widetilde{\delta_{k}^{\rm smp}|}+|\widetilde{\delta^{\rm nst}_{k}}|}{\widetilde{\omega}_{k}}.$ (79)
Now, we examine the asymptotic behavior of the mean absolute error as $M\to\infty$, $T=M^{\gamma}$ (with $\gamma>0$) while $K$ and the resolution of the spatial grid are kept constant. From Eqs. (71) and (72), we have $\widetilde{\omega}_{k}=O(T^{-\beta})$. Then, taking expectation in Eq. (79) with respect to the distribution of the local spectra $\sigma_{\ell}(x)$ (see G) and using Eqs. (74) and (76), we obtain, again up to a constant factor,
@@
$\operatorname{\mathbb{E}}|\delta{\widehat{F}}(\vartheta)|\leq O(T^{\beta}M^{-\nicefrac{{1}}{{2}}})+O(T^{\beta-2})).$ (80)
@@
Here, taking $\beta<2$ and $\gamma<1/(2\beta)$, we have $\operatorname{\mathbb{E}}|\delta{\widehat{F}}(\vartheta)|\to 0$ as $M\to\infty$. Recalling that $f_{\ell}=F(\vartheta(\ell))$, we conclude that the local spectrum $f_{\ell}(x)$ can be estimated from the set of sample band variances $d_{j}(x)$, and the mean absolute error vanishes as the ensemble size goes to infinity. Hence, due to Markov’s inequality, the estimator is consistent provided that the bandpass filters’ transfer functions have the appropriate shape and their widths grow with the sample (ensemble) size.
Appendix J A parametric Locally Stationary Convolution Model
@@
Here we introduce a parametric locally stationary convolution model used in section 5 as a prior model to train the neural spectral estimator and as a model of “truth” to test the LSEF analysis on the sphere. The model generates (in a hierarchical manner) realizations of the spatially variable random local spectra $f_{\ell}(x)$ and realizations of the respective non-stationary random field $\xi(x)$.
J.1 The lowest level in the hierarchy
@@
At the first (i.e., lowest) level in the hierarchy is the random field $\xi(x)$ defined conditionally on the local spectrum $f_{\ell}(x)$ according to Eq. (6), where $\sigma_{\ell}(x)=\sqrt{f_{\ell}(x)}$.
The local spectrum $f_{\ell}(x)$ itself is postulated to obey the following parametric model:
$f_{\ell}(x)=\frac{c(x)}{1+\left[\lambda(x)\ell)\right]^{\gamma(x)}}.$ (81)
@@
Here $c(x)$ is the normalizing variable that ensures that ${\mathsf{Var}}\,\xi(x)=\sum_{\ell}\frac{2\ell+1}{4\pi}f_{\ell}(x)=s^{2}(x)$ and $s(x)$, $\lambda(x)$, $\gamma(x)$ are random fields, which we call parameter fields. They are defined in the next subsection.
J.2 Parameter fields

### v5

@@ Appendix F Bandpass filtering of a locally stationary signal
$\frac{\operatorname{\mathbb{E}}\xi^{2}}{\Lambda^{2}}=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\operatorname{\mathbb{E}}\left(\frac{\partial\sigma_{\ell}}{\partial x}\right)^{2}$ (64)
(cf. Eqs. (12) and (13)). Heuristically, $H(\ell+q)$ in Eq. (63) is then much flatter than $\widetilde{\sigma}_{\ell q}$ (both being considered as functions of $q$). As a result, within the width of $\widetilde{\sigma}_{\ell q}$ (regarded as a function of $q$ with the fixed $\ell$), it holds that $H(\ell+q)\approx H(\ell)$. This suggests the approximation, $\utilde{\varphi}(x)$, that results from replacing, in Eq. (63), $H(\ell+q)$ with $H(\ell)$:
$\varphi(x)\approx\utilde{\varphi}(x)=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\sigma_{\ell}(x)\,\widetilde{\alpha}_{l}\,H(l)\,\mathrm{e}^{\mathsf{i}\ell x}.$ (65)
Its spherical counterpart is given in Eq. (15). Now we study the approximation error
$\delta\varphi(x)=\utilde{\varphi}(x)-\varphi(x)=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\widetilde{\alpha}_{\ell}\,\mathrm{e}^{\mathsf{i}\ell x}\sum_{q=-Q}^{Q}\left[H(\ell)-H(\ell+q)\right]\,\widetilde{\sigma}_{\ell q}\,\mathrm{e}^{\mathsf{i}qx}.$ (66)
It has mean zero and variance
$\operatorname{\mathbb{E}}\left[|\delta\varphi(x)|^{2}{\mbox{$\,|\,$}}\sigma\right]=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\sum_{q=-Q}^{Q}\sum_{q^{\prime}=-Q}^{Q}\left[H(\ell)-H(\ell+q)\right]\left[H(\ell)-H(\ell+q^{\prime})\right]\,\widetilde{\sigma}_{\ell q}\,\widetilde{\sigma}^{*}_{\ell q^{\prime}}\,\mathrm{e}^{\mathsf{i}(q-q^{\prime})x}.$ (67)
We take here expectation over the randomness of the spectral functions $\sigma_{\ell}(x)$. Since the vector-valued random process $\boldsymbol{\sigma}(x)=\{\sigma_{\ell}(x)\}_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}$ is postulated to be stationary, see section 3.6, it holds that $\operatorname{\mathbb{E}}\widetilde{\sigma}_{\ell q}\widetilde{\sigma}_{\ell q^{\prime}}^{*}=0$ for $q\neq q^{\prime}$. As a result, Eq. (67) simplifies to
$\operatorname{\mathbb{E}}|\delta\varphi|^{2}=\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}\left[|\delta\varphi(x)|^{2}\,|\,\sigma\right]=\sum_{\ell=-\ell_{\rm max}}^{\ell_{\rm max}}\sum_{q=-Q}^{Q}\left[H(\ell)-H(\ell+q)\right]^{2}\operatorname{\mathbb{E}}|\widetilde{\sigma}_{\ell q}|^{2}.$ (68)
Now, we note that for each $\ell$, there exists $\bar{\ell}$ such that $H(\ell+q)-H(\ell)=H^{\prime}(\bar{\ell})q=\kappa^{\prime}(\cdot)q/T\equiv\kappa^{\prime}(\cdot)qL_{\cal H}$. Hence
$\operatorname{\mathbb{E}}|\delta\varphi|^{2}=L_{\cal H}^{2}\sum_{\ell}[\kappa^{\prime}(\cdot)]^{2}\sum_{q}q^{2}\operatorname{\mathbb{E}}|\widetilde{\sigma}_{\ell q}|^{2}\leq L_{\cal H}^{2}\sum_{\ell}\operatorname{\mathbb{E}}\left(\frac{\partial\sigma_{\ell}}{\partial x}\right)^{2}=\left(\frac{L_{\cal H}}{\Lambda}\right)^{2}\operatorname{\mathbb{E}}\xi^{2},$ (69)
where Eq. (64) as well as the assumption $|\kappa^{\prime}(\cdot)|\leq 1$ were used.
Note that $\operatorname{\mathbb{E}}|\utilde{\varphi}(x)|^{2}=\operatorname{\mathbb{E}}\operatorname{\mathbb{E}}\left[|\utilde{\varphi}(x)|^{2}\,|\,\sigma\right]=\sum H^{2}(\ell)\operatorname{\mathbb{E}}|\sigma_{\ell}|^{2}$ does not change neither with $\Lambda$ nor with $L_{\cal H}$ as long as the mean process spectrum $\operatorname{\mathbb{E}}\sigma_{\ell}^{2}$ is not changed, as we assume. Therefore the relative approximation error variance tends to zero as $L_{\cal H}/\Lambda\to 0$ at the same rate $(L_{\cal H}/\Lambda)^{2}$ as the absolute error variance.
Finally, we assess the mean absolute approximation error in Eq. (17), $\operatorname{\mathbb{E}}|\delta^{\rm nsta}(x)|$, where $\delta^{\rm nsta}(x)=\utilde{v}(x)-v(x)$, $v(x)=\operatorname{\mathbb{E}}(|\varphi(x)|^{2}{\mbox{$\,|\,$}}\sigma)$, and $\utilde{v}(x)=\operatorname{\mathbb{E}}(|\utilde{\varphi}(x)|^{2}{\mbox{$\,|\,$}}\sigma)$. We have $|\delta^{\rm nsta}|=\bigl{|}\operatorname{\mathbb{E}}\bigl{(}|\utilde{\varphi}|^{2}-|\varphi|^{2}{\mbox{$\,|\,$}}\,\sigma\bigr{)}\bigr{|}\leq\operatorname{\mathbb{E}}\bigl{(}\bigl{|}|\utilde{\varphi}|^{2}-|\varphi|^{2}\bigr{|}\,{\mbox{$\,|\,$}}\sigma\bigr{)}=\operatorname{\mathbb{E}}\bigl{(}\bigl{|}|\utilde{\varphi}|-|\varphi|\bigr{|}(|\utilde{\varphi}|+|\varphi|)\,{\mbox{$\,|\,$}}\sigma\bigr{)}$. Further, the inequality $\bigl{|}|a|-|b|\bigr{|}\leq|a-b|$ entails that $|\delta^{\rm nsta}|\leq\operatorname{\mathbb{E}}\bigl{(}|\delta\varphi|(|\utilde{\varphi}|+|\varphi|)\,{\mbox{$\,|\,$}}\sigma\bigr{)}$. Next, taking expectation with respect to the randomness of the spectral functions $\sigma$, we obtain $\operatorname{\mathbb{E}}|\delta^{\rm nsta}|\leq\operatorname{\mathbb{E}}\bigl{(}|\delta\varphi|(|\utilde{\varphi}|+|\varphi|)\bigr{)}$. Here, the Cauchy-Schwarz inequality yields $\operatorname{\mathbb{E}}|\delta^{\rm nsta}|\leq\sqrt{\operatorname{\mathbb{E}}|\delta\varphi|^{2}}\sqrt{\operatorname{\mathbb{E}}\bigl{(}|\utilde{\varphi}|+|\varphi|\bigr{)}^{2}}$. Finally, assuming that $\operatorname{\mathbb{E}}\bigl{(}|\utilde{\varphi}|+|\varphi|\bigr{)}^{2}$ is bounded above, we conclude from Eq. (69) that the mean absolute non-stationarity error in $\utilde{v}(x)$, that is, the error in the approximate Eq. (17), is
$\operatorname{\mathbb{E}}|\delta^{\rm nsta}|=O\left(\frac{L_{\cal H}}{\Lambda}\right)$ (70)
as ${L_{\cal H}}/{\Lambda}\to 0$.
Appendix G Proof of Proposition 4.1: Consistency of the estimator
Here, we examine the accuracy of the linear estimator described in section 4.4. We consider the large sample limit, $K\to\infty$. The degree of the spatial non-stationarity is kept fixed in this limit. We show that the bandpass filters can be selected in such a way that the estimation error vanishes, implying consistency of the estimator.
We start by recalling that according to Eqs. (17) and (18), the $j$th band variance can be represented as
$v_{j}=\utilde{v}_{j}+\delta^{\rm nsta}_{j},$ (71)
where $\utilde{v}_{j}$ is linearly related to the local spectrum and $\delta^{\rm nsta}_{j}$ is the error of the linear approximation (here and in the sequel, dependencies on the spatial point $x$ are dropped).
Then, following section 4.4, we replace the sum in Eq. (18) with the integral:
@@
where $\vartheta$ is the log-wavenumber, $F(\vartheta)$ represents the local spectrum in the sense that $f_{\ell}=\sigma_{\ell}^{2}=F(\vartheta(\ell))$, $\omega_{j}(\vartheta;T)$ stands for $\frac{2\ell+1}{4\pi}H_{j}^{2}(\ell)$, and $T$ is a real number that controls the widths of the functions $\omega_{j}(\cdot;T)$ (the way it does so is specified below). The discrepancy $\delta^{\rm nsta}_{j}$ is studied in F, where it is generically denoted as $\delta^{\rm nsta}$.
We specify the transfer functions $\omega_{j}(\vartheta,T)$ of the bandpass filters in Eq. (72) as follows. First, we let the number of the filters, $J$, be equal to $M+1$, where $M$ is the bandwidth of the local spectrum, see Definition 4.1. Second, we define the function $\Omega:{\mathbb{R}}\to{\mathbb{R}}$ as a positive even function that is monotonic on ${\mathbb{R}}_{+}$, vanishes at infinity, and satisfies for some $\beta>0$
@@
where $\vartheta_{j}=\frac{\pi}{M}(j-1)$, $j=1,2,\dots,J$, determines the $j$th passband’s central log-wavenumber. We note that by definition, see Eqs. (74) and (75), $T$ determines the length scale of the resulting filters’ impulse response functions: $L_{{\cal H}_{j}}=O(T^{-1})$ as $T\to\infty$. From Eq. (70), we obtain
$\operatorname{\mathbb{E}}|\delta^{\rm nsta}_{j}|=O(T^{-1}).$ (76)
The data we have are the sample band variances $d_{j}$, which estimate the band variances $v_{j}$ up to the sampling error $\delta^{\rm samp}_{j}$:
$d_{j}=v_{j}+\delta^{\rm samp}_{j}.$ (77)
@@
$\operatorname{\mathbb{E}}|\delta^{\rm samp}_{j}|\leq\sqrt{\operatorname{\mathbb{E}}(\delta^{\rm samp}_{j})^{2}}\sim\sqrt{\frac{2}{K}}\operatorname{\mathbb{E}}v_{j}=O(K^{-\nicefrac{{1}}{{2}}})$ (78)
as $K\to\infty$. From Eqs. (71), (72), (75), and (77), we have
$\int_{{\mathbb{S}}^{1}}\omega(\vartheta_{j}-\vartheta,T)\;F(\vartheta)\,{\rm d}\vartheta=d_{j}-\delta^{\rm samp}_{j}-\delta^{\rm nsta}_{j}.$ (79)
@@
This is a linear inverse problem to be solved with respect to $F(\cdot)$ from the data $\{d_{j}\}_{j=1}^{J}$. The left-hand side of Eq. (79) is the convolution $\omega(\cdot;T)*F(\cdot)$ evaluated at $\vartheta_{j}$. This convolution can readily be computed in Fourier space, see Eq. (22), where Eq. (79) becomes
$\widetilde{F}_{m}\,\widetilde{\omega}_{m}(T)\propto\widetilde{d}_{m}-\widetilde{\delta_{m}^{\rm samp}}-\widetilde{\delta^{\rm nsta}_{m}},$ (80)
@@
where the variables with the tilde are Fourier transforms of the respective variables without the tilde and $0\leq m\leq M$. From Eq. (80), we can write down the solution in Fourier space, $\widetilde{\widehat{F}}_{m}\propto{\widetilde{d}_{m}}/{\widetilde{\omega}_{m}(T)}$, and its error (up to a constant factor):
$|\delta\widetilde{\widehat{F}}_{m}|\leq\frac{|\widetilde{\delta_{m}^{\rm samp}|}+|\widetilde{\delta^{\rm nsta}_{m}}|}{\widetilde{\omega}_{m}(T)}.$ (81)
From Eqs. (73) and (74), it follows that $1/\widetilde{\omega}_{m}(T)=O(T^{\beta})$ where, we recall, $\beta>0$. We substitute this latter asymptotic relation in Eq. (81) and take expectation of the resulting equation with respect to the distribution of the random spectral functions $\sigma_{\ell}(\cdot)=\sqrt{f_{\ell}(\cdot)}$ (see section 3.6). We examine the asymptotic behavior of the mean absolute error $\operatorname{\mathbb{E}}|\delta{\widehat{F}}(\vartheta)|$ as $K\to\infty$ and, importantly, $T=K^{\gamma}$ (where $\gamma>0$), with $M={\rm{const}}$. From Eq. (81), we obtain, again up to a constant factor,
@@
$\operatorname{\mathbb{E}}|\delta{\widehat{F}}(\vartheta)|\leq\operatorname{\mathbb{E}}\sum_{m=-M+1}^{M}|\delta\widetilde{\widehat{F}}_{m}|\leq O(T^{\beta}K^{-\nicefrac{{1}}{{2}}})+O(T^{\beta-1}).$ (82)
Here the first inequality is because $\delta{\widehat{F}}(\vartheta)=\sum_{m=-M}^{M}\delta\widetilde{\widehat{F}}_{m}\,\mathrm{e}^{\mathsf{i}m\vartheta}$ and therefore $|\delta{\widehat{F}}(\vartheta)|\leq\sum_{m=-M}^{M}|\delta\widetilde{\widehat{F}}_{m}|$. The second inequality is due to Eqs. (81), (76), (78), and the assumption that $M$ is finite and fixed.
@@
From Eq. (82), taking $10<\beta<$ 1 and $0<\gamma<1/(2\beta)$, we have $\operatorname{\mathbb{E}}|\delta{\widehat{F}}(\vartheta)|\to 0$ as $K\to\infty$. Recalling that $f_{\ell}=F(\vartheta(\ell))$, we conclude that the local spectrum $f_{\ell}(x)$ can be estimated from the set of sample band variances $d_{j}(x)$, and the mean absolute estimation error vanishes as the ensemble size goes to infinity.
It is worth reiterating that this result is obtained in the limit $K\to\infty$ while the degree of non-stationarity is held fixed.
Appendix H The Parametric Locally Stationary Convolution Model
@@
Here we introduce a parametric spatial model used in section 6 as a hyperprior model to train the neural estimator and as a model of ‘truth’ to test the LSEF analysis on the sphere. The model generates (in a hierarchical manner) realizations of the spatially variable random spectral functions $\sigma_{\ell}(x)$ and realizations of the respective non-stationary random field $\xi(x){\mbox{$\,|\,$}}\sigma$.
H.1 General design
@@
At the first (i.e., the lowest) level in the hierarchy is the random field $\xi(x)$ defined conditionally on the spectral functions $\sigma_{\ell}(x)$ according to Eq. (10).
At the second level are the random spectral functions $\sigma_{\ell}(x)$, which are postulated to obey the following parametric model:
$\sigma_{\ell}(x)=\sqrt{\frac{c(x)}{1+\left[\lambda(x)\ell)\right]^{\gamma(x)}}}.$ (83)
@@
Here $c(x)$ is the normalizing variable that ensures that ${\mathsf{Var}}\,\xi(x)=\sum_{\ell}\frac{2\ell+1}{4\pi}\sigma^{2}_{\ell}(x)=s^{2}(x)$ and $s(x)$, $\lambda(x)$, $\gamma(x)$ are random fields, which we call parameter fields.
H.2 Parameter fields